Rebecca is x years old.
Mary is 8 years older than Rebecca.
Jill is three times older than Mary.
The sum of their ages is 67.
Form an equation in terms of x
when you find x work out how old mary is

Answer by Boreal(15235)

Start with Rebecca, the youngest.
R=x
Mary=x+8
Jill is 3(x+8)=3x+24
They add to 67
so 5x+32=67
x=7 (Rebecca)
Mary is is 15